OpenStudy (kirbykirby):

Show that E(X) is equal to this formula, on the interval [a, b]: \[E(X)=a+\int_a^b[1-F(x)]dx\] E(X) is the expectation of X, F(x) is the cdf of X

OpenStudy (kirbykirby):

X is just some random variable. It can be continuous, discrete or mixed

hartnn (hartnn):

CDF = F(x)=P(X \(\le \)x) so, 1-F(x) =P(X \(\ge \)x) =\(\int \limits_x^{\infty }f(x)dx\) you can use this.

OpenStudy (anonymous):

i forget but i think there is some gimmick here where you take \[\int_0^{\infty} xf(x)dx\]and rewrite it as \[-\int_0^{\infty}x(-f(x))dx\] then use parts but i could be wrong

hartnn (hartnn):

now, change the order of double integration....

OpenStudy (kirbykirby):

Oh i think I might have figured it out after all ._. the integration by parts triggered smtg

hartnn (hartnn):

sorry, that should be \(\int \limits_x^{\infty }f(t)dt\) oh...okk.

OpenStudy (kirbykirby):

hm i think double int would work too . thanks !

OpenStudy (zarkon):

double integral is the easy way

hartnn (hartnn):

if you are comfortable with changing the order of integration, it'll give u d answer in few steps...